A chemist has a beaker of a 3% acid solution and a beaker of a 7% acid solution. He needs to make 75ml of a 4% acid solution. complete the table.
a.
3%solution + 7%solution = 4%solution
amount of solution(ml) ____x + ______y =
amount of acid (ml) ____x + ______y = 0.04(75)
b. Use the informaiton in the table to write a system of linear equations.
c. Solve the system of equations to find how much he will use from each beaker.

Found 2 solutions by richwmiller, checkley77:
Answer by richwmiller(17219)

3%solution + 7%solution = 4%solution
total solution 75 ml
3% solution uses x and 7 % uses 75-x
.3x+.7(75-x)=.4(75)

Answer by checkley77(12844)

3%solution + 7%solution = 4%solution
amount of solution(ml) x + y = 75 OR x=75-y
amount of acid (ml) .03x + .07y = 0.04(75)
.03(75-y)+.07y=3
2.25-.03y+.07y=3
.04y=3-2.25
.04y=.75
y=.75/.04
y=18.75 ml of 7% solution is used.
x+18.75=75
x=75-18.75
x=56.25 ml. of 3% solution is used.
Proof:
.03*56.25+.07*18.75=.04*75
1.6875+1.3125=3
3=3